OpenStudy (anonymous):

Graph the two equations to solve the system. y = 2x - 4 2x + 3y = -12

OpenStudy (anonymous):

by soving equations replace 2x in second eqaution with y+4+3y=-12 4y=-12-4 4y=-16 y=-4 substitute y val in 1st equa -4=2x-4 2x=-8 x=-2

OpenStudy (anonymous):